OpenStudy (anonymous):

If 8+ k (in square root form)=15,thenk=

OpenStudy (anonymous):

u have \[\sqrt{8+k}=15\] square both sides what will u get?

OpenStudy (anonymous):

soooo......

OpenStudy (anonymous):

8+k=225 k=217